High-Low Method

Apply the high-low method of cost analysis to the three cost data groups shown below.

Volume
(applicable to
each group)
Group A
Costs
Group B
Costs
Group C
Costs
2,000 $5,200 $2,000 $4,800
12,000 25,200 12,000 4,800
16,000 33,200 16,000 4,800
20,000 41,200 20,000 4,800

What cost behavior patterns are apparent? Express each as a cost formula: Total cost = Fixed costs + (Variable cost x Volume).

Group Fixed costs + (Variable cost x Volume)
A Answer + Answer x Volume
B Answer + Answer x Volume
C Answer + Answer x Volume
Answer #1

variable cost per unit = (41200-5200)/18000 = 2': Fixed cost = 41200-(20000*2) = 1200

Variable cost per unit = (20000-2000)/18000 = 1; Fixed cost = 20000-20000 = 0

What cost behavior patterns are apparent? Express each as a cost formula: Total cost = Fixed costs + (Variable cost x Volume).

Group Fixed costs + (Variable cost x Volume)
A 1200 + 2 *X
B 0 + 1 *X
C 4800 + 0 *X